OpenStudy (anonymous):

what fraction of 8 tons is 1.000lbs?

OpenStudy (anonymous):

There are 2000lbs in a ton. So 8 tons is 16000lbs. YOu probably got it from there.

OpenStudy (anonymous):

\[\frac{ 1 ton }{ 2000 lb } *\frac{ 1 lb }{ 8 ton }\] Simplify this and you should get the fraction i think.

OpenStudy (anonymous):

\[\frac{ 1 }{ 16000 }\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

YES

OpenStudy (mathstudent55):

@hina000 Do you mean one pound or one thousand pounds? If it's 1 lb, then (1 lb)/(8 tons) * (1 ton)/(2000 lb) = 1/16,000 If it's 1000 lb, then (1000 lb)/(8 tons) * (1 ton)/(2000 lb) = 1/16